OpenStudy (diamond77108):

will medal PLZZ HELP!!! What is 4(x−2)=3(2y−1) in standard form?

OpenStudy (retireed):

4(x−2)=3(2y−1) Convert to standard form.... Standard Form: the standard form of a line is in the form Ax + By = C where A is a positive integer, and B, and C are integers. 4x - 8 = 6y - 3 Finish the algebra yourself.

OpenStudy (diamond77108):

i did that but i get 4x+ 6y = 11 and thats not correct

OpenStudy (retireed):

You didn't do the algebra correctly. Try again. Add 8 to both sides and subtract 6y from both sides.

OpenStudy (diamond77108):

ok 4x -6y = 5 thx